# Neural Solver Selection for Combinatorial Optimization

### v1

@@ 3.2 Selection model
We employ a Multiple-Layer Perception (MLP) to predict compatibility scores of neural solvers, where a higher score indicates that it is more suitable to allocate the instance to the corresponding neural solver. This MLP model takes the instance representation and the instance scale $N$ as input and outputs a score vector, where the value of each index is the score of the corresponding neural solver. In summary, the graph encoder and the MLP are cascaded to compose a neural selection model, which can produce the compatibility scores of individual solvers from the raw COP instance in an end-to-end manner. Advanced neural solver features can be incorporated for richer information, as discussed in Section 5. However, we find that even using fixed indices of neural solvers has already been effective, which will be clearly shown in our experiments.
We train the selection model using a supervised dataset comprising thousands of synthetic COP instances. The objective values obtained by the neural solvers are recorded as supervision information. Intuitively, a neural solver with a lower objective value (for minimization) has a higher compatibility score. To learn such desired score output, we employ two losses from the perspectives of classification and ranking.
@@
The selection problem is formulated as a classification task, where the most suitable neural solver for a given instance serves as the ground truth label. By employing classification loss functions such as cross-entropy loss used in our experiments, we can train a selection model. However, this approach focuses on identifying the optimal neural solver and ignores sub-optimal solvers, which may lead to unsatisfactory performance when the selection is inaccurate.
@@
where $g_{\theta}$ denotes the selection model with parameters $\theta$, $I$ denotes a problem instance, and $\phi_{I}$ is the ground-truth ranking on instance $I$. This ranking loss can leverage the dominance relationship of all neural solvers, including sub-optimal ones, which thus can make the selection more robust.
@@ 3.3 Selection strategies
Considering that the intricate structures of COPs may pose great challenge to the selection model, besides greedy selection, we propose several compromise strategies that allow multiple solvers for a single instance based on the confidence level of the selection model, aiming to improve the overall performance with little extra cost.
Greedy selection.
The most straightforward approach is the greedy selection, which chooses the neural solver with the highest score. This method is efficient since only one solver is executed per instance. However, it may be inaccurate, potentially leading to sub-optimal performance.
Top-$k$ selection.
The top-$k$ selection method can be adopted for better optimality, where we select and execute the neural solvers with top-$k$ scores for each instance, thus constructing a portfolio of multiple solvers. This approach increases the likelihood of including the optimal solver but incurs additional computational overhead due to the execution of multiple solvers.
Rejection-based selection.
To balance efficiency and effectiveness, we propose the rejection-based selection strategy, which adaptively selects greedy or top-$k$ selection. Recognizing that the confidence of the greedy selection varies across instances, an advanced strategy is to employ the top-$k$ selection for low-confidence instances to enhance performance and utilize only the greedy selection for high-confidence ones to minimize computational cost. To implement this strategy, we can use a confidence measure to determine whether to accept or reject the greedy selection. If the confidence in the greedy selection is below a threshold, we reject it and apply the top-$k$ selection to the instance for improved optimality. In this paper, we adopt the simple yet effective softmax response (Hendrycks & Gimpel, 2017) as the confidence measure, and define the threshold by rejecting a certain fraction of test instances with the lowest confidence levels.
Top-$p$ selection.
We further propose a top-$p$ selection strategy that selects the smallest subset of solvers whose normalized scores (i.e., selection probabilities) sum up to at least $p$. The value of $p$ is predefined or adjusted according to the time budget. Thus, this strategy adaptively determines the number of selected neural solvers by covering a certain amount of probability mass, rather than relying on a fixed number $k$.
@@ 4 Experiments
To examine the effectiveness of our proposed selection framework, we conduct experiments on TSP and CVRP, investigating the following Research Questions (RQ): RQ1: How does the proposed selection framework perform compared to individual neural solvers? RQ2: How does the proposed selection framework perform when the problem distribution shifts and the problem scale increases? RQ3: How do different implementations of components affect the performance of the framework? We introduce the experimental settings in Section 4.1 and investigate the above RQs in Section 4.2. The code and data used in our experiments are provided in the supplementary materials.
@@ 4.1 Experimental settings
Synthetic TSP and CVRP.
We generate synthetic TSP and CVRP instances by sampling node coordinates from Gaussian mixture distributions with randomized covariance matrices. In the case of CVRP, vehicle capacities are generated using either the scale-related capacity or the triangular distribution. We consider varying problem scales, where the scale $N$ is sampled uniformly from $[50,500]$. More details of the data generation process are provided in Appendix A.2.
Datasets.
@@
For training, we generate $10,000$ TSP and CVRP instances and apply $8$-fold instance augmentation (Kwon et al., 2020). For test, we generate smaller synthetic datasets comprising $1,000$ instances. Figures 1(a) and 1(b) in Section 1 are based on results from the CVRP test dataset. To evaluate the out-of-distribution performance, we utilize two well-known benchmarks with more complex problem distributions and larger problem scales (up to $N=1002$): TSPLIB (Reinelt, 1991) and CVRPLIB Set-X (Uchoa et al., 2017). For TSPLIB, we select a subset of instances with $N\leq 1002$, and CVRPLIB Set-X includes instances ranging from $N=100$ to $1000$. These problem scales are larger than the scale $N\in[50,500]$ of our training datasets.
Open-source neural solvers.
@@
We choose recent open-source neural solvers with state-of-the-art performance as the candidates, including Omni (Zhou et al., 2023), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), DIFUSCO (Sun & Yang, 2023), T2T (Li et al., 2023), ELG (Gao et al., 2024), INViT (Fang et al., 2024) and MVMoE (Zhou et al., 2024b). Greedy decoding is used for all the methods to avoid stochasticity. We set the pomo size to $100$ and the augmentation number to $8$ for the methods based on POMO (Kwon et al., 2020). The number of denoising steps is set to $50$ and the number of $2$-opt iterations is set to $100$ for diffusion-based methods. These individual solvers constitute a neural solver zoo. Ideally, if we can always select the best solver from the zoo for each instance, the optimal performance is achieved, which is also the performance upper bound of our selection model. Considering that some neural solvers contribute little to the overall performance, we iteratively eliminate the least contributive solver from the candidates, resulting in a more compact neural solver zoo. This process reduces the zoo size to $7$ solvers for TSP and $5$ for CVRP. Further details of the elimination procedure are provided in Appendix A.3.
Hyperparameters.
(1) Hyperparameters of graph encoders. For the graph attention encoder, we set the number of layers to $4$. For the hierarchical graph encoder, we use $2$ blocks where each block has $2$ attention layers. The embedding dimension is set to $128$. Other hyperparameters of encoders can be found in Appendix A.1. (2) Hyperparameters of training. The Adam optimizer (Kingma & Ba, 2015) is employed for training, where we set the learning rate to $1\times 10^{-4}$ and the weight decay to $1\times 10^{-6}$. The number of epochs is set to $50$. The final model is chosen according to the performance on a validation dataset with $1,000$ synthetic instances. We train 5 selection models using different random seeds and report the mean and standard deviation of their performance. (3) Hyperparameters of selection strategies. For the top-$k$ strategy, we set $k=2$. For the rejection-based strategy, we reject the 20% of instances with the lowest confidence levels (i.e., the highest selection probability of all individual solvers), and apply top-$2$ selection to these rejected instances. For the top-$p$ strategy, we set $p=0.5$ for TSP and $p=0.8$ for CVRP.
Performance metrics.
@@
Following previous studies, we employ the gap to the best-known solution $\frac{c_{I}(\hat{\sigma})-c_{I}(\sigma^{*})}{c_{I}(\sigma^{*})}$ as the performance metric, called optimality gap, where $\hat{\sigma}$ is the solution obtained by each method, $\sigma^{*}$ is the best-known solution computed by extensive search of expert solvers (Helsgaun, 2017; Vidal, 2022), and $c_{I}()$ is the cost function of problem instance $I$. We also report the average time to evaluate efficiency, which includes both the running of neural solvers and selection.
@@ 4.2 Experimental results
RQ1: How does the proposed selection framework perform compared to individual neural solvers?
In Table 1, we present the performance of several implementations of our selection framework on synthetic TSP and CVRP, alongside the results of the top-$3$ individual neural solvers***DIFUSCO and T2T have multiple trained models. We only report the best results of these models. . We can observe that all implementations of our framework outperform the best neural solver on both TSP and CVRP, demonstrating the effectiveness of our framework. For example, using ranking loss and the top-$k$ selection strategy with $k=2$, our framework achieves average optimality gaps of 1.51% on TSP and 4.82% on CVRP, surpassing the best individual solver’s gaps of 2.33% on TSP and 6.82% on CVRP, achieved by DIFUSCO and Omni, respectively. Moreover, except utilizing the top-$k$ strategy, our selection framework is nearly as efficient as running a single solver. In some cases, our framework can obtain better optimality gaps while consuming even less time. For instance, using ranking loss and greedy selection on TSP leads to the average optimality gap 1.86% with 1.33s, while the best individual solver DIFUSCO achieves 2.33% gap with 1.45s. In Table 1, OPT (the fourth row) denotes the optimal performance for selection, which is obtained by running all individual solvers for each instance and selecting the best one. The best optimality gaps achieved by our selection framework (using ranking loss and top-$k$ selection with $k=2$) are close to OPT, with gaps of 1.51% on TSP and 4.81% on CVRP, compared to OPT’s gaps of 1.24% on TSP and 4.64% on CVRP. Furthermore, our framework can offer significant speed advantages over OPT, e.g., consuming an average time of 2.56s on TSP, whereas OPT requires an average time of 8.93s. Note that complete results for all individual solvers are provided in Appendix A.10.
Table 1: Empirical results on synthetic TSP and CVRP datasets, reporting the mean (standard deviation) over five independent runs. The top three individual solvers are included for comparison, and OPT denotes the optimal performance for selection, which is computed by running all individual solvers in the zoo for each instance and selecting the best one. The best individual solver and its results are underlined, and the best optimality gaps, excluding OPT, are highlighted in boldface.
@@
Extension of RQ1: Is the performance of the top-$k$ selection better than the solver portfolio with the same size?
The top-$k$ strategy enhances the performance by running a selected subset of the solver zoo for each instance, which certainly costs more time than individual solvers. For a fair comparison, we benchmark our top-$k$ selection method against a solver portfolio of the same size $k$. We construct this solver portfolio by exhaustively enumerating all possible subsets of size $k$ and selecting the one with the best overall performance. As shown in Appendix A.5, our top-$k$ selection consistently outperforms the size-$k$ solver portfolio across $k=\{1,2,3,4\}$ on all datasets, i.e., TSP, CVRP, TSPLIB and CVRPLIB Set-X, demonstrating the effectiveness of our selection model.
RQ2: How does the proposed selection framework perform when the problem distribution shifts and the problem scale increases?
We evaluate the generalization performance on two benchmarks, TSPLIB and CVRPLIB Set-X, which contain out-of-distribution and larger-scale instances. As shown in Table 2, all implementations of our selection framework generalize well, where the ranking model using top-$k$ selection improves the optimality gap by 0.88% (i.e., 1.95%-1.07%) on TSPLIB and by 0.71% (i.e., 6.10%-5.39%) on CVRPLIB Set-X, compared to the best individual solvers T2T and ELG on these two benchmarks. These results show that our selection framework is robust against the distribution shifts and increases in problem scale.
RQ3: How do different implementations affect performance?
We evaluate and compare different implementations of the three components in our framework:
@@
(1) Feature extraction methods. We compare the manual features (Smith-Miles et al., 2010) (see Appendix A.4), graph attention encoder Kool et al. (2019), and hierarchical graph encoder in Table 3. All methods are trained using ranking loss, and we report the optimality gap with greedy selection. As shown in Table 3, even the simplest manual features perform well, achieving better results than the best individual solver across three datasets — TSP, CVRP, and TSPLIB. This further validates the effectiveness of our selection framework. Comparing the third and fourth columns, we observe that the graph attention encoder consistently outperforms manual features on all datasets, verifying the superiority of learned features. Furthermore, by comparing the fourth and fifth columns, we find that while the graph attention encoder has already been effective on synthetic datasets, introducing the hierarchical encoder can further improve generalization performance on out-of-distribution datasets, TSPLIB and CVRPLIB Set-X, which is quite important in practice. This enhanced generalization capability may be attributed to the hierarchical encoder’s ability to leverage the inherent hierarchical structures in COPs. More ablation studies of the hierarchical encoder are provided in Appendix A.6.
@@
(2) Loss functions to train the selection model. We can clearly observe from Tables 1 and 2 that the model trained with ranking loss generally outperforms the one trained with classification loss, particularly when employing top-$p$ selection or under out-of-distribution settings. We also compare their accuracy of selecting the best single individual, which is similar as shown Appendix A.8. Thus, the benefit of ranking loss over classification loss shows the importance of incorporating the dominance relationships among sub-optimal solvers, which can make the selection more robust.
@@
(3) Selection strategies. Greedy selection is efficient by selecting only the predicted best solver. Instead, top-$k$ selection selects the best $k$ solvers for better optimality gaps, but resulting in longer time. Rejection-based and top-$p$ selection provide a trade-off between optimality gap and time. Here, we focus on the evaluation of rejection-based and top-$p$ selection. We tune their hyperparameters (e.g., rejection ratio, $k$, and $p$) to obtain a range of results, provided in Appendix A.7. The results show the rejection-based selection with smaller $k$ ($k=2$ or $3$) tends to achieve better trade-off. Comparing top-$p$ selection and rejection-based selection, their performance has no significant difference. This is expected, because both of their principles are running more individual neural solvers when the confidence of the selection model is insufficient. However, the top-$p$ selection may be preferable in practice, where only one hyperparameter $p$ is associated.

### v2

@@ 3.2 Selection model
We employ a Multiple-Layer Perception (MLP) to predict the compatibility scores of neural solvers, where a higher score indicates that it is more suitable to allocate the instance to the corresponding neural solver. This MLP model takes the instance representation and the instance scale $N$ as input and outputs a score vector, where the value of each index is the score of the corresponding neural solver. In summary, the graph encoder and the MLP are cascaded to compose a selection model, which can produce the compatibility scores of solvers for the COP instances in an end-to-end manner. Advanced neural solver features can be incorporated for richer information, as discussed in Section 5. However, we find that even using fixed indices of neural solvers can be effective, which will be clearly shown in our experiments. We train the selection model using a supervised dataset comprising thousands of synthetic COP instances. The objective values obtained by the neural solvers are recorded as supervision information. Intuitively, a neural solver with a lower objective value (for minimization) has a higher compatibility score. To learn it, we employ two losses from the perspectives of classification and ranking.
@@
The selection problem is formulated as classification, where the most suitable neural solver for a given instance serves as the ground truth label. By employing classification loss functions such as cross-entropy loss in our experiments, we can train a selection model. However, this approach focuses on identifying the optimal neural solver and ignores sub-optimal solvers, which may lead to unsatisfactory performance when the selection is inaccurate.
@@
where $g_{\theta}$ denotes the selection model with parameters $\theta$, $I$ denotes a problem instance, and $\phi_{I}$ is the ground-truth ranking on instance $I$. This ranking loss can leverage the dominance relationship of all neural solvers, including sub-optimal ones, which can make the selection more robust.
@@ 3.3 Selection strategies
Considering that the intricate structures of COPs may pose great challenges to the selection model, besides greedy selection (choosing the highest-scoring solver), we propose several compromise strategies that allow multiple solvers for a single instance based on the confidence level of the selection model, aiming to improve the overall performance with little extra cost.
  • Top-$k$ selection. The top-$k$ selection method can be adopted for better optimality, where we select and execute the neural solvers with top-$k$ scores for each instance, thus constructing a portfolio of multiple solvers. This approach increases the likelihood of including the optimal solver but incurs additional computational overhead due to the execution of multiple solvers.
  • Rejection-based selection. To balance efficiency and effectiveness, we propose the rejection-based selection strategy, which adaptively selects greedy or top-$k$ selection. Recognizing that the confidence of greedy selection varies between instances, an advanced strategy is to employ the top-$k$ selection for low-confidence instances to enhance performance and utilize only greedy selection for high-confidence ones to minimize computational cost. To implement this strategy, we can use a confidence measure to determine whether to accept or reject greedy selection. If the confidence in greedy selection is below a threshold, we reject it and apply the top-$k$ selection to the instance. In this paper, we adopt the simple yet effective softmax response (Hendrycks & Gimpel, 2017) as the confidence measure and define the threshold by rejecting a certain fraction of test instances with the lowest confidence levels.
  • Top-$p$ selection. We further propose a top-$p$ selection strategy that selects the smallest subset of solvers with normalized scores sum up to at least $p$. It can adaptively determine the number of selected neural solvers by covering a certain amount of confidence, rather than relying on a fixed number $k$. The value of $p$ is predefined or adjusted according to time budget.
@@ 4 Experiments
To examine the effectiveness of our proposed selection framework, we conduct experiments on TSP and CVRP, investigating the following Research Questions (RQ): RQ1: How does the proposed selection framework perform compared to individual neural solvers? RQ2: How does the proposed selection framework perform when the problem distribution shifts and the problem scale increases? RQ3: How do different implementations of components affect the performance of the framework? We introduce the experimental settings in Section 4.1 and investigate the above RQs in Section 4.2.
@@ 4.1 Experimental settings
We generate synthetic TSP and CVRP instances by sampling node coordinates from Gaussian mixture distributions (more details of the data generation are provided in Appendix A.2.1). For training, we generate $10,000$ TSP and CVRP instances and apply $8$-fold instance augmentation (Kwon et al., 2020). For test, a smaller synthetic datasets with $1,000$ instances is used. Figures 1(a) and 1(b) in Section 1 depict results on the CVRP test dataset. To evaluate the out-of-distribution performance, we utilize two well-known benchmarks with more complex problem distributions and larger problem scales (up to $N=1002$): TSPLIB (Reinelt, 1991) and CVRPLIB Set-X (Uchoa et al., 2017). For TSPLIB, we select a subset of instances with $N\leq 1002$, and CVRPLIB Set-X includes instances ranging from $N=100$ to $1000$. These problem scales are larger than $N\in[50,500]$ in our training datasets. Notably, in Appendix A.10, we also include experiments on larger scales up to $N=2000$ for further evaluation.
@@
Open-source neural solvers. We choose recent open-source neural solvers with state-of-the-art performance as the candidates, including Omni (Zhou et al., 2023), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), DIFUSCO (Sun & Yang, 2023), T2T (Li et al., 2023), ELG (Gao et al., 2024), INViT (Fang et al., 2024) and MVMoE (Zhou et al., 2024b). More details about the realization can be found in Appendix A.2.2. Considering that some neural solvers contribute little to the overall performance, we iteratively eliminate the least contributive solver from the candidates, resulting in a more compact neural solver zoo. This process reduces the zoo size to $7$ solvers for TSP and $5$ for CVRP. Further details of the elimination procedure are provided in Appendix A.3.
Hyperparameters. Limited by space, we provide the details of hyperparameters in our experiments in Appendix A.2.3, including the hyperparameters of graph encoders, hyperparameters of training, and hyperparameter of selection strategies.
@@
Performance metrics. Following previous studies, we use the gap to the best-known solution $\frac{c_{I}(\hat{\sigma})-c_{I}(\sigma^{*})}{c_{I}(\sigma^{*})}$ as the performance metric, called optimality gap, where $\hat{\sigma}$ is the solution obtained by each method, $\sigma^{*}$ is the best-known solution computed by extensive search of expert solvers (Helsgaun, 2017; Vidal, 2022), and $c_{I}()$ is the cost function of problem instance $I$. We also report the average time (including both the running of solvers and selection) to evaluate efficiency, .
@@ 4.2 Experimental results
RQ1: How does the proposed selection framework perform compared to individual neural solvers? In Table 1, we present the performance of several implementations of our selection framework on synthetic TSP and CVRP, alongside the results of the top-$3$ individual neural solvers [^1] DIFUSCO and T2T have multiple trained models. We only report the best results of these models. . We can observe that all implementations of our framework outperform the best neural solver on both TSP and CVRP, demonstrating the effectiveness of our framework. For example, using ranking loss and the top-$k$ selection strategy with $k=2$, our framework achieves average optimality gaps of 1.51% on TSP and 4.82% on CVRP, surpassing the best individual solver’s gaps of 2.33% on TSP and 6.82% on CVRP, achieved by DIFUSCO and Omni, respectively. Moreover, except utilizing the top-$k$ strategy, our selection framework is nearly as efficient as running a single solver. In some cases, our framework can obtain better optimality gaps while consuming even less time. For instance, using ranking loss and greedy selection on TSP leads to the average optimality gap 1.86% with 1.33s, while the best individual solver DIFUSCO achieves 2.33% gap with 1.45s. In Table 1, Oracle (the fourth row) denotes the optimal performance for selection, which is obtained by running all individual solvers for each instance and selecting the best one. The best optimality gaps achieved by our selection framework (using ranking loss and top-$k$ selection with $k=2$) are close to Oracle, with gaps of 1.51% on TSP and 4.81% on CVRP, compared to Oracle’s gaps of 1.24% on TSP and 4.64% on CVRP. Furthermore, our framework can offer significant speed advantages over Oracle, e.g., consuming an average time of 2.56s on TSP, whereas Oracle requires an average time of 8.93s. Note that complete results for all individual solvers are provided in Appendix A.12.
Table 1: Empirical results on synthetic TSP and CVRP datasets, reporting the mean (standard deviation) over five independent runs. The top three individual solvers are included for comparison, and Oracle denotes the optimal performance for selection, which is computed by running all individual solvers in the zoo for each instance and selecting the best one. The best individual solver and its results are underlined, and the best optimality gaps, excluding Oracle, are highlighted in boldface.
@@
Extension of RQ1: Is the performance of the top-$k$ selection better than the solver portfolio with the same size? The top-$k$ strategy enhances the performance by running a selected subset of the solver zoo for each instance, which certainly costs more time than individual solvers. For a fair comparison, we benchmark our top-$k$ selection method against a solver portfolio of the same size $k$. We construct this solver portfolio by exhaustively enumerating all possible subsets of size $k$ and selecting the one with the best overall performance. As shown in Appendix A.6, our top-$k$ selection consistently outperforms the size-$k$ solver portfolio across $k=\{1,2,3,4\}$ on all datasets, i.e., TSP, CVRP, TSPLIB and CVRPLIB Set-X, demonstrating the effectiveness of our selection model.
RQ2: How does the proposed selection framework perform when the problem distribution shifts and the problem scale increases? We evaluate the generalization performance on two benchmarks, TSPLIB and CVRPLIB Set-X, which contain out-of-distribution and larger-scale instances. As shown in Table 2, all implementations of our selection framework generalize well, where the ranking model using top-$k$ selection improves the optimality gap by 0.88% (i.e., 1.95%-1.07%) on TSPLIB and by 0.71% (i.e., 6.10%-5.39%) on CVRPLIB Set-X, compared to the best individual solvers T2T and ELG on these two benchmarks. These results show that our selection framework is robust against the distribution shifts and increases in problem scale.
RQ3: How do different implementations affect performance? We evaluate and compare different implementations of the three components in our framework:
@@
(1) Feature extraction methods. We compare the manual features (Smith-Miles et al., 2010) (see Appendix A.4), graph attention encoder (Kool et al., 2019), and hierarchical graph encoder in Table 3. All methods are trained using ranking loss, and we report the optimality gap with greedy selection. As shown in Table 3, even the simplest manual features perform well, achieving better results than the best individual solver across three datasets — TSP, CVRP, and TSPLIB. This further validates the effectiveness of our selection framework. Comparing the third and fourth columns, we observe that the graph attention encoder consistently outperforms manual features on all datasets, verifying the superiority of learned features. More manual features are compared in Appendix A.5. Furthermore, by comparing the fourth and fifth columns, we find that while the graph attention encoder has already been effective on synthetic datasets, introducing the hierarchical encoder can further improve generalization performance on out-of-distribution datasets, TSPLIB and CVRPLIB Set-X, which is quite important in practice. This enhanced generalization capability may be attributed to the hierarchical encoder’s ability to leverage the inherent hierarchical structures in COPs. More ablation studies of the hierarchical encoder are in Appendix A.7.
@@
(2) Loss functions to train the selection model. We can clearly observe from Tables 1 and 2 that the model trained with ranking loss generally outperforms that trained with classification loss, particularly when employing top-$p$ selection or under out-of-distribution settings. We also compare their accuracy of selecting the best single individual, which is similar as shown Appendix A.9. Thus, the benefit of ranking loss over classification loss shows the importance of incorporating the dominance relationships among sub-optimal solvers.
@@
(3) Selection strategies. Greedy selection is efficient by selecting only the predicted best solver. Instead, top-$k$ selection selects the best $k$ solvers for better optimality gaps, but resulting in longer time. Rejection-based and top-$p$ selection provide a trade-off between optimality gap and time. Here, we focus on the evaluation of rejection-based and top-$p$ selection. We tune their hyperparameters (e.g., rejection ratio, $k$, and $p$) to obtain a range of results, provided in Appendix A.8. The results show that the rejection-based selection with smaller $k$ ($k=2$ or $3$) tends to achieve better trade-off. Comparing top-$p$ and rejection-based selection, their performance has no obvious difference. This is expected as their principles are both running more individual neural solvers when the confidence of the selection model is insufficient. However, the top-$p$ selection may be preferable in practice since only one hyperparameter $p$ is associated.
4.3 Exploration on neural solver features
To enable generalization to unseen neural solvers, we propose a preliminary feature extraction method that leverages representative instances to characterize each neural solver. Specifically, for a given neural solver, we sort those instances where the neural solver performs the best in ascending order according to the ratio of the objective value that the solver obtains to the runner-up objective value and select the top 1% as its representative instances. Then, we use an instance encoder to obtain embeddings for representative instances as their token vectors. A two-layer transformer model further processes these token vectors to learn a summarized feature, which serves as the feature representation of the neural solver.
To integrate a newly added neural solver, we first identify its representative instances from a synthetic dataset and employ the aforementioned networks to compute its feature representation. The selection model can then leverage the newly added solver by considering its feature during selection, without the need for any fine-tuning. The implementation details of each component are described as follows.
Instance tokenization.
We use a hierarchical graph encoder as the tokenization encoder to generate embeddings for each representative instance. To stabilize the instance tokens during training, we update $\theta^{\prime}$ using a momentum-based moving average of the parameters $\theta$ of the instance feature encoder: $\theta^{\prime}\leftarrow m\cdot\theta^{\prime}+(1-m)\cdot\theta$, where $m\in[0,1)$ is a momentum coefficient (We set $m=0.99$ in experiments). Only the parameters $\theta$ are updated via back-propagation. This momentum update ensures that $\theta^{\prime}$ evolves more smoothly than $\theta$, resulting in stable instance tokenization.
Transformer architecture.
Selection model with neural solver features.
The selection model integrates both the instance features and the neural solver features to output a score for each instance-solver pair. We employ an MLP to compute these scores. For each instance, the scores across all neural solvers are normalized to derive the probability distribution.
To evaluate the effectiveness of this method, we remove the second-best neural solver from the current solver zoo, train the selection model using the supervision information over the pruned solver zoo, and reintroduce the removed solver during testing. Figure 3 presents the top-$k$ selection performance with and without the newly added (extra) solver. The results show that the performance with the newly added solver is generally better than the performance without it, demonstrating that the selection model can leverage the information of unseen solvers without any finetuning. In other words, the selection model can generalize to unseen solvers. However, we observe a slight decrease in top-$1$ performance, indicating that the current feature design is not accurate enough for the most precise top-$1$ selection. This highlights the need for further refinement and optimization to our preliminary feature extraction method.
Figure 3: Performance of introducing an extra neural solver based on the neural solver feature. The results of top-$k$ selection with and without extra neural solver are presented. The horizontal axis represents the number of selected solvers.